OpenStudy (vshiroky):

Suppose the growth of population during the period from 1980 to 2008 can be approximated by an equation of the form y=mx+b, where x is the number of years since 1980 and y is the population measured in millions. Find the equation if 2 ordered pairs that satisfy it are (0,234) and (10,255)

OpenStudy (vshiroky):

I think it's y=21/10x+234

OpenStudy (anonymous):

m=change in y/change in x=(255-234)/(10-0)=21/10, b=y intercept= 234, therefore y=(21/10)x+234, you are right.